(d) Pre-emptive priority scheduling [4 marks]:

Consider the following set of processes, with the length of the CPU-burst time given in milliseconds:

Process

Burst

Priority

Arrival Time

P1

8

3

0

P2

3

4

1

P3

6

2

3

P4

3

1

5

P5

1

5

7

P6

3

8

14

P7

8

5

18

(d-1) Draw Gantt chart illustrating the execution of these processes using pre-emptive priority (a smaller priority number implies a higher priority) [2 marks]

00   01   02   03   04   05   06   07   08   09   10   11   12   13   14   15   16   17   18   19   20   21   22   23   24   25   26   27   28   29   30   31   32   33   34

(d-2) Compute the turnaround time of all processes in the previous question [2 mark]

Process

Turnaround Time

P1

P2

P3

P4

P5

P6

P7

Average Turnaround Time
Answer #1

Q(d)

(d)-1

Sol:

We have the required Gantt Chart as follows:

Gantt Chart P1 P3 P4 P3 P1 P2 P5 P7 P6 3 5 8 12 17 20 21 29 32

Which can be explained as follows:

  • P1 enters at time=0 so P1 is executed first.
  • P2 enters at t=1 but its priority is less than P1 so P1 continues execution
  • P3 enters at t=3 with higher priority than P1 so P3 starts execution and P1 is pre-empted
  • Now, P3 does execution till t=5.
  • P4 enters execution at t=5 since it now has highest priority and it will complete its execution till t=8 because of highest priority.
  • Now, after P4, P3 has the highest priority so it will complete its execution till t=12
  • Now, after P3, P1 has the highest priority so it will complete its execution till t=17
  • Now, after P1, P2 has the highest priority so it will complete its execution till t=20
  • Now, after P2, P5 has the highest priority so it will complete its execution till t=21
  • Now, after P5, P7 has the highest priority so it will complete its execution till t=29
  • Now, after P5, P6 will complete its execution till t=32

(Note that all times are in millisecond(ms))

(d)-2

Sol:

We know that Turnaround Time = Completion Time - Arrival Time

So, we have the required table as follows (here completion time is found using the gantt chart where last executing time of the process is written):

Process Arrival Time(in ms) Completion Time(in ms) Turnaround Time(in ms)
P1 0 17 17 - 0 = 17
P2 1 20 20 - 1 = 19
P3 3 12 12 - 3 = 9
P4 5 8 8 - 5 = 3
P5 7 21 21 - 7 = 14
P6 14 32 32 - 14 = 18
P7 18 29 29 - 18 = 11

Average Turnaround Time = Sum of Turnaround times of all the processes/Number of processes

= (17 + 19 + 9 + 3 + 14 + 18 + 11)/7

= 91/7

= 13 ms (Ans.)

(Thank You !!, Please consider upvoting if it was helpful !!)